OpenStudy (anonymous):

HELP FOR A MEDAL!! ms.Campbell invested $7200 for 2 1/2 years. If she earned $756, at what rate(R) did she invest her money?

Vocaloid (vocaloid):

money earned = (investment)*(1+rate/100)^time

Vocaloid (vocaloid):

plug in your numbers and solve for r

OpenStudy (anonymous):

I did ....7200=(756) (?) (2 1/2) I=PRT

OpenStudy (anonymous):

I multiplied (756) * (2 1/2) and got 1890 and then divide 7200/ 1890 and got 3.8095238 it's wrong :(

Vocaloid (vocaloid):

756 = (7200)(1+rate/100)^(2.5)

Vocaloid (vocaloid):

(756/7200) = (1+rate/100)^2.5

Vocaloid (vocaloid):

(756/7200)^(1/2.5) = 1 + rate/100

Vocaloid (vocaloid):

(756/7200)^(1/2.5) - 1 = rate/100

OpenStudy (anonymous):

I got 0.042

Vocaloid (vocaloid):

oh, whoops, small mistake on my part money earned = 7200+756 ([7200+756]/7200)^(1/2.5) - 1 = rate/100 solve for rate

OpenStudy (anonymous):

Hmmm would it be cross multiply 1.105 1/2.5 ?? :)